### v1

# Focus: Efficient Keyframe Selection for Long Video Understanding

Abstract
Multimodal large language models (MLLMs) represent images and video frames as visual tokens. Scaling from single images to hour-long videos, however, inflates the token budget far beyond practical limits. Popular pipelines therefore either uniformly subsample or apply keyframe selection with retrieval-style scoring using smaller vision-language models. However, these keyframe selection methods still rely on pre-filtering before selection to reduce the inference cost and can miss the most informative moments.
We propose Focus, Frame-Optimistic Confidence Upper-bound Selection, a training-free, model-agnostic keyframe selection module that selects query-relevant frames under a strict token budget. Focus formulates keyframe selection as a combinatorial pure-exploration (CPE) problem in multi-armed bandits: it treats short temporal clips as arms, and uses empirical means and Bernstein confidence radius to identify informative regions while preserving exploration of uncertain areas. The resulting two-stage exploration-exploitation procedure reduces from a sequential policy with theoretical guarantees, first identifying high-value temporal regions, then selecting top-scoring frames within each region On two long-video question-answering benchmarks, Focus delivers substantial accuracy improvements while processing less than 2% of video frames. For videos longer than 20 minutes, it achieves an 11.9% gain in accuracy on LongVideoBench, demonstrating its effectiveness as a keyframe selection method and providing a simple and general solution for scalable long-video understanding with MLLMs. Code is available at https://github.com/NUS-HPC-AI-Lab/FOCUS.
1 Introduction
“The art of being wise is the art of knowing what to overlook.” — William James
Recent advances in large language models (LLMs) and multimodal LLMs (MLLMs) have significantly improved visual understanding and reasoning. In current frameworks, images are encoded into visual tokens aligned with text and jointly processed by the LLM. Extending this paradigm to videos—especially long, untrimmed ones—introduces a key challenge: the sheer number of frames leads to an overwhelming number of visual tokens, making inference computationally prohibitive.
A common solution is aggressive downsampling (Wang et al., 2022b; Lin et al., 2023; Maaz et al., 2024; Zhang et al., 2025c), but uniformly sampling a handful of frames (e.g., 64 from a one-hour video) often misses critical content (Tang et al., 2025; Zhang et al., 2025b). Increasing the frame rate, on the other hand, causes token explosion (Wang et al., 2024c). This trade-off motivates the need for keyframe selection: choosing a small set of informative frames that preserve semantics while staying within token limits.
Recent methods address this by scoring frame relevance with pre-trained vision-language encoders (e.g., CLIP (Radford et al., 2021) or BLIP (Li et al., 2022)) and then pick the highest-relevance frames (Tang et al., 2025; Zhang et al., 2025b). These text–image matching approaches are typically training-free and plug in easily before the visual encoder in MLLM stacks, retrieving frames with higher relevance other than uniform sampling. Despite their success, current keyframe selection methods still face scalability and efficiency limitations. For a one-hour video at 30 fps (over $10^{5}$ frames), exhaustively scoring all frames entails on the order of $10^{11}$-$10^{12}$ FLOPs with a vision-language encoder like BLIP (Li et al., 2022). This scaling pressure forces existing methods to uniformly sample the video to lower frame rate before the scoring process. This pre-filtering process before keyframe selection undermines the goal of identifying most informative keyframes from all frames (Zhang et al., 2025b; Tang et al., 2025).
Figure 1: Temporal autocorrelation (ACF) of per-frame query relevance on LongVideoBench and Video-MME. We compute frame-level relevance per video and take the ACF over time lags (seconds); solid lines show the median across videos and shaded bands the interquartile range. The dashed line marks the correlation half-life level ($\rho(\delta)=0.5$).
In this work, we propose Focus, Frame-Optimal Confidence Upper-Bound Selection, a training-free, plug-and-play keyframe selection method designed to process extremely long videos with minimal computational overhead. Focus is easy to implement in practice while offering an elegant theoretical foundation.
The key insight behind Focus is grounded in the observation that natural videos exhibit strong temporal locality: adjacent frames are highly correlated in appearance and motion (Wiegand et al., 2003; Wang et al., 2016; 2022b). This local smoothness naturally extends to frame–query relevance scores. As illustrated in Figure 1, we compute the autocorrelation function (ACF) of relevance scores $r_{t}$ on LongVideoBench and VideoMME. The results show a strong local correlation structure, with a half-life of approximately 5 seconds. This observation implies that exhaustive scoring of all frames is unnecessary. Instead, we can formulate keyframe selection as a bandit problem to adaptively allocate computation: quickly filtering out irrelevant temporal regions, concentrating scoring on promising segments, and ultimately prioritizing the most informative keyframes.
Focus first partitions the video into short temporal clips, each treated as an arm in a multi-armed bandit. The clip selection is then framed as a Combinatorial Pure-Exploration (CPE) problem: the goal is to identify a subset of arms that maximizes expected cumulative relevance under a strict token budget. Each arm maintains an empirical mean relevance and a Bernstein-style confidence radius. Computation is adaptively allocated to clips that are either promising (high mean) or uncertain (large confidence radius), following an optimism-in-the-face-of-uncertainty principle. This iterative process enjoys theoretical convergence guarantees. To leverage parallel computation without sacrificing optimism, we reduce the iterative strategy to a coarse-to-fine schedule: optimistic means guide exploration, while unbiased empirical means inform final arm selection. Within each selected arm, we extract the top-relevance frames to construct the final keyframe set.
We validate the effectiveness of our approach on two video understanding benchmarks, including LongVideoBench (Wu et al., 2024) and Video-MME (Fu et al., 2025). The proposed Focus is tested as an off-the-shelf module on with four popular MLLMs. Focus improves answer accuracy over state-of-the-art keyframe selection baselines across benchmarks while maintaining lower inference cost. The gains are especially pronounced on long-form videos: for videos longer than 20 minutes on LongVideoBench, Focus delivers a 1.9% accuracy improvement while still cutting inference cost.
In summary, our main contributions are three-fold: (1) We formulate query-aware keyframe selection as a budgeted combinatorial pure-exploration (CPE) problem in a multi-armed bandit setting; (2) We introduce Focus, a training-free, model-agnostic keyframe selection module that selects query-relevant frames under a strict token budget; (3) We validate the effectiveness of Focus on two long-video understanding benchmarks, achieving consistent gains across four popular MLLMs.
2 Method
Figure 2: Overview of Focus. Focus partitions videos into fixed-length clips as bandit arms, applies optimistic confidence upper-bound arm selection and selects final keyframes within each promising arms.
2.1 Problem Formulation
Keyframe Selection Setup.
Let a video be $V=({\bm{x}}_{1},\ldots,{\bm{x}}_{T})$ and denote the corresponding text query as $q$. Let the frame index set be ${\mathbb{T}}=\{1,\ldots,T\}$. A downstream multimodal LLM $\Phi$ consumes a subset of frames indexed by ${\mathbb{K}}\subseteq{\mathbb{T}}$ with $|{\mathbb{K}}|=k$ and produces an answer $\hat{a}=\Phi\big(q,\{{\bm{x}}_{t}\}_{t\in{\mathbb{K}}}\big)$. Let $R_{\Phi}({\mathbb{K}}\mid V,q)$ denote the task-level utility of the selected frames (e.g., quality of generated answer, relevance to query, or other performance metrics).
Oracle and Surrogate Objective.
The oracle objective chooses $\mathbb{K}$} to maximize expected utility:
${\mathbb{K}}^{\mathrm{oracle}}(V,q)=\operatorname*{arg\,max}_{{\mathbb{K}}\subseteq{\mathbb{T}},\,|{\mathbb{K}}|=k}\;\mathbb{E}\big[R_{\Phi}({\mathbb{K}}\mid V,q)\big],$ (1)
Direct optimization to equation 1 is infeasible due to the combinatorial search space and the high cost of black-box evaluations of $\Phi$. We further expand the task-level utility $R_{\Phi}({\mathbb{K}}\mid V,q)$ to a summation of frame-level utility $y_{t}\in[0,1]$:
${\mathbb{K}}^{\star}=\operatorname*{arg\,max}_{{\mathbb{K}}\subseteq{\mathbb{T}},\,|{\mathbb{K}}|=k}\;\mathbb{E}\big[\sum_{t\in{\mathbb{K}}}y_{t}\big].$ (2)
However, estimating the contribution of each frame $t$ to the task-level utility is also intractable. We therefore posit that $y_{t}$ is indirectly observable via a vision-language encoder $\psi$ that outputs a relevance score $r_{t}=\psi({\bm{x}}_{t},q;{\bm{\theta}})=y_{t}+\epsilon_{\psi}$, where $\epsilon_{\psi}$ denotes encoder-induced noise. We assume $\epsilon_{\psi}$ follows some distribution that are supported on $[0,1]$ and with zero mean and $\sigma_{\psi}^{2}$ variance. Under this assumption, the relevance score $r_{t}$ is a unbiased estimator of $y_{t}$ which is also commonly used in many works (Tang et al., 2025; Yu et al., 2024) implicitly.
Exhaustively scoring all $T$ frames to get $\{r_{t}\}$ is computationally prohibitive, especially for hourly long videos which contains over $10^{5}$ frames. This computational constraint motivates us to model keyframe selection under budget constraints, where we strategically allocate a limited sampling budget to identify the most promising temporal segments before producing the final set of $k$ keyframes. Instead of directly optimizing equation 2 at the frame level, we will approximate it through a combinatorial pure-exploration multi-armed bandit formulation at the clip level, which significantly reduces exploration cost.
2.2 Clip-Level Selection as Multi-Armed Bandit
For a video $V=({\bm{x}}_{1},\ldots,{\bm{x}}_{T})$, we partition the timeline into $M$ non-overlapping fixed-length clips $\mathcal{A}=\{A_{a}\}_{a=1}^{M}$, where each clip $A_{a}\subseteq{\mathbb{T}}$ spans frames $s_{a},e_{a}]$ and is treated as a bandit arm. We assume that frame-level utility within the same arm share the same distribution: $y_{t}\sim\mathcal{\nu}_{a}$ for all $t\in[s_{a},e_{a}]$, where $\mathcal{\nu}_{a}$ has mean $\mu_{a}$ and variance $\sigma_{a}^{2}$. We define pulling the arm $a$ as randomly sampling one frame from that clip and observing its query relevance score $r_{t}$ as a reward.
Intuitively, our goal is to focus on the most promising clips which means we have to identify the optimal subset $S^{\star}\subseteq\mathcal{A}$. Formally, we define the decision class ${\mathbb{S}}\in 2^{\mathcal{A}}$ as a subset of the power set of $\mathcal{A}$. The optimal member $S^{\star}$ of decision class $\mathbb{S}$} is defined as
$S^{\star}=\operatorname*{arg\,max}_{S\in{\mathbb{S}}}\;\sum_{a\in S}\mu_{a}.$ (3)
Under the classic CPE framework, the learner’s objective is to identify $S^{\star}$ after interacting with the arms over a sequence of rounds. In the keyframe selection setting, our final goal is to further select $k$ keyframes from the selected arms. Denote $\{k_{a}\}_{a=1}^{|S^{\star}|}$ as the number of keyframes allocated to the $a$-th selected arm. We further define the frame-level optimal keyframe subset ${\mathbb{K}}_{a}^{\star}$ as
${\mathbb{K}}_{a}^{\star}=\operatorname*{arg\,max}_{{\mathbb{K}}_{a}\subseteq A_{a},\,|{\mathbb{K}}_{a}|=k_{a}}\;\sum_{t\in{\mathbb{K}}_{a}}y_{t}.$ (4)
The final keyframe subset $\mathbb{K}$$^{\star}$ is then defined as ${\mathbb{K}}^{\star}=\bigcup_{a\in S^{\star}}{\mathbb{K}}_{a}^{\star}$. Empirically, we assume the decision class $\mathbb{S}$} is all size-$m$ subsets of $\mathcal{A}$ and keyframes are equally distributed across the promising arms. This setting gives us an elegant theoretical guarantee of regret bound as shown in section C and is also proved to be effective in our experiments.
2.3 Optimistic Confidence Upper-bound Arm Selection
Algorithm 1 Iterative Optimistic Confidence Upper-bound Arm Selection
1:Maximization oracle $\mathrm{TopM}(\{\mu_{a}\},m)\to{\mathbb{A}}\subseteq\mathcal{A}$
2:Initialize: Empirical means $\hat{\mu}_{0}(a)\leftarrow 0$ and $N_{0}(a)\leftarrow 0$ for all $a$.
3:Pull each arm $a\in\mathcal{A}$ for $q$ times and observe the rewards.
4:Update empirical means $\hat{\mu}_{a}$ for all $a$.
[…]
6:for $n\leftarrow mq,mq\!+\!1,\ldots$ do
7:  ${\mathbb{A}}_{n}\leftarrow\mathrm{TopM}(\hat{{\bm{\mu}}},m)$
8:  Compute confidence radius $\beta_{a}(n)$ for all $a\in\mathcal{A}$ $\triangleright$ $\beta_{a}(n)$ defined in equation 5
9:  for $a\leftarrow 1$ to $M$ do
10:   if $a\in{\mathbb{A}}_{n}$ then
11:     $\tilde{\mu}_{n}(a)\leftarrow\hat{\mu}_{n}(a)-\beta_{a}(n)$
12:   else
13:     $\tilde{\mu}_{n}(a)\leftarrow\hat{\mu}_{n}(a)+\beta_{a}(n)$
14:   end if
15:  end for
16:  $\tilde{{\mathbb{A}}}_{n}\leftarrow\mathrm{TopM}(\tilde{{\bm{\mu}}},m)$
17:  if $\tilde{{\mathbb{A}}}_{n}={\mathbb{A}}_{n}$ then
18:   return ${\mathbb{A}}_{n}$
19:  end if
20:  $p_{n}\leftarrow\operatorname*{arg\,max}\limits_{a\in(\tilde{{\mathbb{A}}}_{n}\setminus{\mathbb{A}}_{n})\,\cup\,({\mathbb{A}}_{n}\setminus\tilde{{\mathbb{A}}}_{n})}\beta_{a}(n)$ $\triangleright$ break ties arbitrarily
21:  Pull arm $p_{n}$ and observe the reward
22:  Update empirical means $\hat{\mu}({p_{n}})$ with the observed reward
23:  $N_{n+1}(p_{n})\leftarrow N_{n}(p_{n})+1$
24:end for
Algorithm 2 Optimistic Confidence Upper-bound Arm Selection
1:Maximization oracle $\mathrm{TopM}(\{\mu_{a}\},m)\to{\mathbb{A}}\subseteq\mathcal{A}$
2:Initialize: Empirical means $\hat{\mu}_{0}(a)\leftarrow 0$ and $N_{0}(a)\leftarrow 0$ for all $a$.
3:// Stage I: Coarse exploration
4:Pull each arm $a\in\mathcal{A}$ for $q$ times and observe the rewards.
5:Update empirical means $\hat{\mu}_{a}$ for all $a$.
[…]
7:Compute confidence radius $\beta_{a}(n)$ for all $a\in\mathcal{A}$
[…]
10:// Stage II: Fine-grained exploitation
11:Pull each arm $a\in{\mathbb{A}}_{\text{coarse}}$ for $z$ times and observe the rewards.
12:Update empirical means $\hat{\mu}_{a}$ for $a\in{\mathbb{A}}_{\text{coarse}}$
[…]
14:return ${\mathbb{A}}_{\text{fine}}$
2.3.1 Optimal Arm Selection.
Generally, we play a exploration game by pulling an arm $a$ and observing the reward $r_{t}$ at each round $n$. We maintain two core empirical statistics for each arm $a$ during this process: an empirical mean $\hat{\mu}_{a}$ and an empirical Bernstein confidence radius (variance-adaptive) $\beta_{a}$, following the UCV-V style bound (Audibert et al., 2009):
$\beta_{a}(n)\;=\;\sqrt{\frac{2\,\hat{\sigma}_{a}^{2}\,\ln n}{\max(1,N_{a}(n))}}\;+\frac{3\,\ln n}{\max(1,N_{a}(n))}.$ (5)
Here $N_{a}(n)$ is the number of pulls for arm $a$ at round $n$ and $n=\sum_{a\in\mathcal{A}}N_{a}(n)$ is the total number of pulls. The confidence radius ensures that the empirical mean is within the confidence radius of the true mean with high probability, i.e.,
$\mathcal{P}\left[\left|\hat{\mu}_{a}-\mu_{a}\right|\leq\beta_{a}(n)\right]\geq 1-\frac{6}{n}.$ (6)
Please refer to Appendix B for the detailed proof.
As shown in Algorithm 1, the optimistic confidence upper-bound arm selection starts with an initialization phase where we pull each arm for $q$ times and observe the relevance scores as rewards. We then update the empirical means $\hat{\mu}_{a}$ and compute the confidence radius $\beta_{a}(n)$ for each arm $a$. Note the relevance score $r_{t}$ is an unbiased estimator of $y_{t}$ so we have $\mathbb{E}[\hat{\mu}_{a}]=\mu_{a}$. Then we choose the best $m$ arms using the empirical means $\hat{\mu}_{a}$, i.e., ${\mathbb{A}}_{n}=\mathrm{TopM}(\hat{{\bm{\mu}}},m)$, where $\hat{{\bm{\mu}}}$ is the vector of all arms’ empirical means and $\mathrm{TopM}(\cdot,m)$ returns a set of the $m$ arms with the largest empirical means.
We further refine the arm selection by evaluating the "potential" of each arm. To be specific, for arm $a\in{\mathbb{A}}_{n}$, we compute the lower confidence bound of the empirical mean, i.e., $\mathrm{LCB}_{a}(n)=\hat{\mu}_{a}-\beta_{a}(n)$; for arm $a\notin{\mathbb{A}}_{n}$, we compute the upper confidence bound of the empirical mean, i.e., $\mathrm{UCB}_{a}(n)=\hat{\mu}_{a}+\beta_{a}(n)$. If
$\max_{a\notin{\mathbb{A}}_{n}}\mathrm{UCB}_{a}(n)\geq\min_{a\in{\mathbb{A}}_{n}}\mathrm{LCB}_{a}(n),$ (7)
this indicates that some arms outside the current top-$m$ set are still potential to be included in the top-$m$ set. Thus, we choose the arm $a$ that we are most uncertain about, i.e.,
$a=\operatorname*{arg\,max}\limits_{a\in(\tilde{{\mathbb{A}}}_{n}\setminus{\mathbb{A}}_{n})\,\cup\,({\mathbb{A}}_{n}\setminus\tilde{{\mathbb{A}}}_{n})}\beta_{a}(n).$ (8)
We then pull this arm $a$ for $q$ times and repeat the process until the top-$m$ set is unchanged, i.e., ${\mathbb{A}}_{n+1}={\mathbb{A}}_{n}$. We then return the top-$m$ set ${\mathbb{A}}_{n}$.
It is easy to see Algorithm 1 is guaranteed to return the optimal top-$m$ set ${\mathbb{A}}_{n}$ with high probability (see Section C for the detailed proof). However, the iterative process is empirically inefficient (or intractable) as the sequential arm-pulls and updating can not be parallelizable. We have to pull the arms one-by-one which means forward the vision-language model with batch size 1 sequentially. This costs significant waste of GPU utilization.
2.3.2 Two-stage Arm Selection.
To make the procedure practical and easy to parallelize, we specialize Algorithm 1 into the two-stage, batch variant in Algorithm 2. The overall framework is shown in Figure 2.
Stage I: Coarse initialization.
We pull each arm $q$ times in parallel and update the empirical means $\hat{\mu}_{a}$ and confidence radii $\beta_{a}(n)$ for all $a\in\mathcal{A}$. This stage coincides with the initialization phase of Algorithm 1 and serves as a coarse exploration pass that produces reliable per-arm statistics at low coordination cost.
Stage II: Fine-grained exploration (batched).
Using the optimistic scores $\tilde{\mu}a=\hat{\mu}a+\beta_{a}(n)$, we select the top $\alpha m$ arms, $\mathcal{A}_{\text{coarse}}=\mathrm{TopM}(\tilde{{\bm{\mu}}},,\alpha m)$, and allocate an additional $z$ pulls to each $a\in\mathcal{A}_{\text{coarse}}$ (performed in a single batch). Here, $\alpha$ is a hyperparameter that controls the ratio of the coarse exploration budget to the fine-grained exploration budget. This stage is a batched counterpart of the iterative loop in Algorithm 1: it implements the “optimism in the face of uncertainty” principle by concentrating samples on arms with the largest UCB values, while avoiding per-step scheduling overhead.
Final Arm Selection.
After the fine exploitation, we form the final set by selecting the best $m$ arms according to the unbiased empirical means, ${\mathbb{A}}_{\text{fine}}=\mathrm{TopM}(\hat{{\bm{\mu}}},m)$. This choice mirrors $\delta$-PAC identification routines, where optimistic scores guide exploration but the recommendation itself is based on $\hat{\mu}_{a}$ rather than $\tilde{\mu}_{a}$.
2.4 Frame Selection within Selected Arms
Given the selected arm set ${\mathbb{A}}_{\text{fine}}$ and a total budget of $K$ frames, we sample $k_{a}$ frames per arm $a\in{\mathbb{A}}_{\text{fine}}$ with equal allocation (i.e., $k_{a}=\mathrm{round}(k/|{\mathbb{A}}_{\text{fine}}|)$, adjusted to sum to $K$). For each arm $a$ with index set ${\mathbb{T}}_{a}$ and observed rewards $\{r_{a,s}\}_{s\in S_{a}}$ at sampled indices $T_{a}\subseteq{\mathbb{T}}_{a}$, we simply interpolate all rewards $\hat{r}_{a,t}$ within the arm using the nearest-neighbor assignment. We then form a per-arm sampling distribution according to the interpolated rewards and draw $k_{a}$ frames without replacement from $p_{a}$. The final keyframe set is $\mathcal{K}=\bigcup_{a\in\mathcal{A}_{\text{fine}}}\mathcal{K}_{a}$.
3 Experiments
3.1 Experimental Setup
Benchmarks
We follow the LMMs-Eval framework Zhang et al. (2024a) and adopt the open-source evaluation protocol from AKS for benchmarks, prompts, and scoring. Our experiments focus on two long-video multiple-choice QA benchmarks: LongVideoBench Wu et al. (2024) and VideoMME Fu et al. (2025). These datasets feature videos lasting up to an hour, where effective keyframe selection becomes crucial for performance. To ensure fair comparison (Tang et al., 2025), we disable subtitles, perform zero-shot evaluation, and keep model parameters frozen—varying only the frame selection strategy (our method versus uniform sampling).
Implementation Details
We test both open-source video MLLMs (Qwen2VL (Wang et al., 2024a), LLaVA-OV (Li et al., 2025), LLaVA-Video (Zhang et al., 2025c) and Qwen2-7B (Yang et al., 2024) language model) and the commercial GPT-4o (0513). For frame relevance scoring, we use BLIP ITM (Li et al., 2022) to compute $r_{t}=\psi({\bm{x}}_{t},q;{\bm{\theta}})$, where $r_{t}$ estimates the latent frame-level utility as described in Section 2.1, which is justified as a promising choice by Tang et al. (2025).
3.2 Performance Analysis
Model #Frame LLM LongVideoBench Video-MME
GPT-4V 256 61.3 59.9
Gemini-1.5-Flash 256 61.6 70.3
Gemini-1.5-Pro 256 64.0 75.0
VideoLLaVA 8 7B 39.1 39.9
MiniCPM-V 2.6 64 8B 54.9 60.9
InternVL2-40B 16 40B 59.7 61.2
LLaVA-Video-72B 64 72B 63.9 70.6
GPT-4o 32 51.6 61.8
GPT-4o w/ Ours 32 54.8 $\uparrow 3.2$ 62.5 $\uparrow 0.7$
Qwen2-VL-7B 32 7B 55.6 57.6
Qwen2-VL-7B w/ Ours 32 7B 62.3 $\uparrow 6.7$ 59.7 $\uparrow 2.1$
LLaVA-OV-7B 32 7B 54.8 56.5
LLaVA-OV-7B w/ Ours 32 7B 60.7 $\uparrow 5.9$ 58.3 $\uparrow 1.8$
LLaVA-Video-7B 64 7B 58.9 64.4
LLaVA-Video-7B w/ Ours 64 7B 63.5 $\uparrow 4.6$ 65.4 $\uparrow 1.0$
Table 1: Video-question answering accuracy (%) of various MLLMs on LongVideoBench and Video-MME. Focus is integrated into GPT-4o, Qwen2-VL, LLaVA-OV, and LLaVA-Video. The suffix “w/ Ours” denotes models using keyframes selected by our method; otherwise, frames are uniformly sampled. #Frame indicates the number of frames provided to the MLLM, and LLM denotes the language model size. We also include performance of additional popular MLLMs for reference.
We evaluate Focus by using it to select keyframes as the visual input for the four aforementioned MLLMs, and compare it against the commonly used uniform sampling strategy. The results on LongVideoBench and Video-MME are summarized in Table 1.
Improved Performance via Frame Selection.
As shown in Table 1, Focus consistently outperforms uniform sampling across both open-source and closed-source MLLMs on both LongVideoBench and Video-MME.
Specifically, on LongVideoBench, Focus improves accuracy by 3.2% on GPT-4o, 6.7% on Qwen2-VL-7B, 5.9% on LLaVA-OV-7B, and 4.6% on LLaVA-Video-7B. On Video-MME, the gains are 0.7%, 2.1%, 1.8%, and 1.0% on the same models, respectively.
[…]
Interpretability through Visualizations.
We visualize the frames selected by Focus alongside uniformly sampled frames for two examples from LongVideoBench and Video-MME in Figure 3.
Note that LongVideoBench and Video-MME differ substantially in how their video-question pairs are constructed. In general, LongVideoBench features more detailed and specific questions, while Video-MME focuses on concise, high-level queries. Moreover, LongVideoBench tends to ask about specific scenes or events, whereas Video-MME emphasizes global understanding of the video content.
To highlight this distinction, we manually mark the most informative frames relative to the query using yellow stars. These frames are more temporally concentrated in LongVideoBench (around specific events) and more uniformly distributed across the timeline in Video-MME.
This difference helps explain why Focus achieves greater performance gains on LongVideoBench: our method assumes that frame-level relevance scores are i.i.d., a common setting in multi-armed bandit formulations. This assumption neglects temporal dependencies between video segments. Consequently, retrieval-based methods for keyframe selection typically require regularization (Tang et al., 2025; Yu et al., 2024) to promote diversity and ensure coverage.
If temporal dependencies between segments (arms) are taken into account, the problem setting shifts toward Lipschitz or metric bandits (Kleinberg et al., 2008; Bubeck et al., 2011), and contextual bandits (Chu et al., 2011; Agarwal et al., 2014). We leave such extensions to future work.
Figure 3: Comparison between uniformly sampled frames and those selected by Focus. The left column shows two examples from LongVideoBench; the right column shows two from Video-MME. Yellow stars indicate manually annotated frames that are most informative to the query, many of which are successfully captured by Focus.
3.3 Comparison with State-of-the-Art
Method LongVideoBench Video-MME
Short Medium Long Overall Short Medium Long Overall
Uniform 67.5 57.4 51.8 58.9 76.4 62.6 54.3 64.4
Top-$K$ 72.3 58.0 60.5 62.3 75.4 60.4 53.0 62.9
AKS 72.3 59.2 56.1 62.1 76.3 62.8 54.7 64.6
Focus (ours) 72.3 59.0 63.7 63.5 76.5 63.5 56.1 65.4
Table 2: Comparison between our method and state-of-the-art keyframe selection baselines under matched keyframe count. Results are reported by video length buckets: Short, Medium, and Long. For Video-MME, we adopt its original categorization: Short (<2 min), Medium (4-15 min), and Long (30-60 min). For LongVideoBench, we define Short as videos shorter than 3 minutes, Medium as 3-20 minutes, and Long as over 20 minutes to ensure a balanced distribution.
To further validate the effectiveness of Focus, we compare it against state-of-the-art training-free keyframe selection methods on both LongVideoBench and Video-MME. Specifically, we consider recent approaches based on vision-language similarity:
  • Top-$K$: Computes relevance scores between each frame and the query, then selects the top-$K$ scoring frames. Due to computational constraints, we apply a pre-filtering step by downsampling videos to 1 frame per second.
  • AKS (Tang et al., 2025): A recent method that adaptively balances frame relevance and temporal coverage. It is considered the current state-of-the-art and also incorporates pre-filtering via downsampling to 1 frame per second (Tang et al., 2025).
Fair comparison protocol. We ensure a fair comparison by: (i) evaluating all methods using LLaVA-Video-7B, the best-performing MLLM in our setup; (ii) fixing the number of selected keyframes to $k=64$; (iii) using the same vision-language encoder (e.g., BLIP) for frame scoring whenever possible. Results are summarized in Table 2.
Consistency across different lengths.
Focus achieves consistent performance gains across all video length categories, with particularly strong improvements on long videos. On LongVideoBench, Focus outperforms uniform sampling by 11.9% and Top-$K$ by 7.6% on videos longer than 20 minutes. On Video-MME, the respective improvements are 1.8% and 1.4%.
We also observe that on short videos, all keyframe selection methods perform similarly and consistently outperform uniform sampling. We attribute this to a possible saturation in the reasoning capabilities of the underlying MLLM (LLaVA-Video-7B), where input selection plays a less critical role.
Efficiency comparison.
We report the efficiency of each method in Table 3, measuring both the number of frames “seen” (i.e., scored by a vision-language model) and the total GPU hours required to perform keyframe selection. All GPU hours are measured using a single NVIDIA H100 (80GB) GPU on the LongVideoBench dataset.
As shown, AKS without pre-filtering is computationally infeasible in practice, as it requires scoring all video frames—amounting to over 255 GPU hours by the optimistic estimation. With pre-filtering, the cost drops significantly to 9.3 GPU hours. In contrast, Focus is the most efficient: it requires only 1.6% of the BLIP forward passes and just 5.5 GPU hours, while simultaneously achieving the best overall performance.
Method Filtering-free Frames Seen (%) GPU hours
AKS w/o pre-filtering
100 255
AKS w/ pre-filtering
3.7 9.3
Focus (Ours)
1.6 5.5
Table 3: Efficiency comparison of keyframe selection methods on LongVideoBench. "Pre-filtering" refers to downsampling videos to 1 fps prior to selection. Note that the official AKS pipeline includes this pre-filtering step by default. “Frames Seen (%)” counts the proportion of frame-level BLIP forward passes relative to scoring all frames; GPU hours are measured on a single H100 (80GB).
3.4 Efficiency-Accuracy Trade-off
Focus exposes a natural trade-off between accuracy and computational cost through a single hyperparameter $\alpha$, which controls the fraction of arms selected for fine-grained exploration. We report accuracy and efficiency under different $\alpha$ settings in Table 4.
Accuracy (%) Frames Seen (%) GPU hours
$\alpha=0.1$ 62.9 1.1 3.5
$\alpha=0.25$ 63.5 1.6 5.5
$\alpha=0.5$ 63.6 2.5 9.2
Table 4: Effect of $\alpha$ on the performance and efficiency of Focus. “Frames Seen (%)” counts the proportion of frame-level BLIP forward passes relative to scoring all frames; GPU hours are measured on a single H100 (80GB).
We observe that choice of $\alpha$ has a significant impact on the efficiency while remain stable on the performance. When $\alpha=0.1$, Focus requires around 1.1% of the frames BLIP forward passes while only 3.5 GPU hours. When $\alpha=0.5$, Focus requires around 2.5% of the frames BLIP forward passes while only 9.2 GPU hours. Exhaustively exploiting all arms would require 9.3 GPU hours, while the performance gain compared to $\alpha=0.25$ is negligible.
4 Conclusion
We addressed the core bottleneck of long-video understanding in MLLMs—the explosion of visual tokens—by introducing Focus, a training-free, plug-and-play keyframe selection method that allocates computation under a strict budget. Focus first partitions the video into temporal clips, treats each as an arm in a bandit problem, and then identifies query-relevant regions via a combinatorial pure-exploration strategy using empirical means and Bernstein confidence bounds. To improve efficiency, we reduce the iterative bandit process to a coarse-to-fine two-stage procedure that preserves optimism while enabling parallel inference.
Experiments on two challenging long-video QA benchmarks demonstrate that Focus consistently improves accuracy across four MLLMs while processing fewer than 2% of video frames. Our results show that lightweight, training-free keyframe selection—when guided by statistical principles—can significantly enhance the scalability and practicality of MLLMs for long-video understanding.
5 Reproducibility Statement
We provide a comprehensive theoretical analysis of our method in Appendix B and Appendix C. The source code for this work is publicly available at https://github.com/NUS-HPC-AI-Lab/FOCUS. All models and datasets used in our study are publicly accessible.
Appendix A Appendix
A.1 Related Work
A.1.1 Multimodal Large Language Models (MLLMs) for Video Understanding
Recent MLLMs extend large language models with visual encoders, encoding images or frames into visual tokens that are fused with text to support open-ended video understanding. Most follow an encode-project-fuse pipeline with instruction tuning, as exemplified by the LLaVA family, Video-LLaVA/Video-LLaMA/Video-ChatGPT, and LLaMA-Vid/VideoChat (Liu et al., 2023; Lin et al., 2023; Zhang et al., 2023; Maaz et al., 2024; Li et al., 2024c; b). Progress has largely come from scaling data/backbones and strengthening cross-modal alignment (MiniCPM-V, InternVL/InternVL2, Qwen2-VL; data-centric and modality-binding advances via ShareGPT4Video and LanguageBind) (Yao et al., 2024; Chen et al., 2024e; d; c; Wang et al., 2024a; Chen et al., 2024a; Zhu et al., 2024), together with architectural refinements that unify multi-granularity visual inputs and tighten temporal adapters, and that improve projector efficiency or curricula (LLaVA-OneVision, LLaVA-NeXT/LLaVA-NeXT-Video, Aria, PLLaVA, Kangaroo) (Li et al., 2025; Liu et al., 2024a; Zhang et al., 2024c; Li et al., 2024a; Xu et al., 2024; Liu et al., 2024b). Finally, several models explicitly target extended context and hierarchical summarization for long-form understanding (LongVILA, LongVA, LongVLM, LongVU) (Chen et al., 2024b; Zhang et al., 2024b; Weng et al., 2024; Shen et al., 2024).
However, this tokenization-first paradigm encounters token explosion on long videos, where dense sampling yields prohibitive sequences. Recent efforts reduce the budget by compressing or restructuring tokens: MovieChat (Song et al., 2024) compacts frames into sparse memory, Video-XL-2 (Qin et al., 2025) synthesizes condensed tokens, and VideoStreaming (Qian et al., 2024) processes streams incrementally to cap tokens. Planning/tool-augmented agents (e.g., VideoAgent (Wang et al., 2024b)) curb perception via selective analysis, while hierarchical controllers (VideoTree (Wang et al., 2025)) and scaling recipes (VideoLLaMA 3 (Zhang et al., 2025a)) aid long-horizon reasoning. Beyond compression, ViLAMP (Cheng et al., 2025) uses mixed-precision tokenization to emphasize differential frames/patches and allocate capacity adaptively; long-context instruction-tuning such as Long-VITA (Shen et al., 2025) complements these strategies for long videos.
A.1.2 Vision-Language Pretrained Models
Cross-modal vision-language pretraining spans two-stream fusion, single-stream fusion, dual-encoder contrastive learning, and encoder-decoder hybrids. Two-stream models such as ViLBERT (Lu et al., 2019) and LXMERT (Tan & Bansal, 2019) encode vision and text separately and fuse via cross-attention, while single-stream counterparts—VisualBERT (Li et al., 2019), VL-BERT (Su et al., 2020), UNITER (Chen et al., 2020)—concatenate region features with text in a unified Transformer using MLM and alignment losses. Large-scale dual encoders like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) learn contrastive embeddings for zero-shot transfer, with FILIP (Yao et al., 2022) improving fine-grained patch-token alignment. Hybrid objectives combine contrastive and generative training (Li et al., 2021; Yu et al., 2022; Wang et al., 2022a; Chen et al., 2023) unify captioning and VQA. The BLIP family integrates vision encoders with language modeling—BLIP (Li et al., 2022) and BLIP-2 (Li et al., 2023) (via a lightweight Q-Former)—while Flamingo (Alayrac et al., 2022) and PaLM-E (Driess et al., 2023) inject visual inputs into large LMs for few-shot multimodal reasoning.
Extending to video, early pretraining models learned joint spatio-temporal-language representations with lightweight fusion and sparse sampling. VideoBERT (Sun et al., 2019) pairs frame sequences with transcripts in a BERT-style objective for retrieval and script generation, while HERO (Li et al., 2020) and ClipBERT (Lei et al., 2021) improve efficiency via hierarchical encoding and key-frame sampling for video-text retrieval and QA. Building directly on large image-text models, Clip4Clip (Luo et al., 2022) reuses CLIP encoders and matches videos to text via contrastive similarity, and FrozenBiLM (Yang et al., 2022) freezes a bi-directional LM while aligning a video encoder for zero-shot VQA.
A.1.3 Keyframe Selection
In video representation learning, keyframe selection spans two major paradigms.
Training-free keyframe selection.
Recent training-free methods leverage pretrained vision-language models and lightweight heuristics to pick informative, query-relevant frames. Adaptive Keyframe Sampling (AKS) maximizes prompt-frame similarity while enforcing temporal coverage via a split-and-judge policy (Tang et al., 2025); Q-Frame ranks frames by query-conditioned importance and preserves a few at higher resolution for detail (Zhang et al., 2025b). Text-frame alignment with frozen models further enables plug-and-play selectors (KeyVideoLLM, BOLT) that boost Video-LLM performance without fine-tuning (Liang et al., 2024; Liu et al., 2025). To avoid redundancy and preserve structure under a token budget, Logic-in-Frames performs dynamic, logic-verified search (Guo et al., 2025), while VideoTree builds a hierarchical, query-adaptive frame pyramid that expands salient scenes (Wang et al., 2025).
Instruction-aligned and learned selectors.
Instruction-guided approaches train selectors with LLM/MLLM feedback: Frame-Voyager learns to query frame combinations by ranking sets with a pretrained Video-LLM (Yu et al., 2024), and Hu et al. (2025) supervise a lightweight selector using MLLM-derived single-frame relevance and multi-frame complementarity. Classical summarization remains relevant: supervised LSTM-based models (vsLSTM, dppLSTM; hierarchical RNNs) learn importance/diversity from human summaries (Zhang et al., 2016; 2018; Zhao et al., 2017), while unsupervised RL/adversarial methods (DR-DSN, SUM-GAN) optimize diversity-representativeness or realism without labels (Zhou et al., 2018; Mahasseni et al., 2017); however, these are typically task-agnostic and may miss frames critical for query-driven VQA.
A.1.4 Multi-Armed Bandits and Batched Exploration
Multi-armed bandits (MAB) encompass both regret minimization and pure exploration. Regret-oriented methods such as UCB variants and Thompson Sampling establish logarithmic-regret foundations for sequential decision-making (Auer et al., 2002; Lai & Robbins, 1985; Agrawal & Goyal, 2012). Pure exploration instead targets high-confidence identification with minimal samples, formalized as best-arm (and top-$k$) identification (Even-Dar et al., 2006; Bubeck et al., 2009; Kalyanakrishnan & Stone, 2010; Cao et al., 2015). Early elimination schemes (Successive/Median Elimination) provide PAC guarantees (Even-Dar et al., 2006; 2002), while confidence-bound and racing families—LUCB, UCB-E, and near-optimal lil’UCB—sharpen sample complexity and approach known lower bounds (Kalyanakrishnan et al., 2012; Audibert & Bubeck, 2010; Karnin et al., 2013; Jamieson et al., 2014; Kaufmann et al., 2016). Beyond single arms, combinatorial pure exploration (CPE) seeks an optimal subset under structural constraints, combining bandit confidence bounds with combinatorial oracles to search exponentially large spaces efficiently (Chen et al., 2016; Lattimore & Szepesvári, 2020).
Fully sequential adaptivity can be impractical when decisions must be made in few rounds or in parallel. Batched (parallel) bandits address this by operating over a small number of adaptivity rounds, yet retain near-sequential sample efficiency for pure exploration in theory and practice (Perchet et al., 2016; Jun et al., 2016; Gao et al., 2019). Batch-elimination/LUCB-style procedures match sequential complexity up to constants with only a handful of updates (Jun et al., 2016), and lower-bound trade-offs between batches and samples are well understood with matching algorithms (Perchet et al., 2016; Kaufmann et al., 2016; Tuynman & Degenne, 2025). Recent designs such as Tri-BBAI attain asymptotically optimal fixed-confidence BAI with just three batches, underscoring the feasibility of resource-constrained exploration (Jin et al., 2024).
Appendix B Bernstein Confidence Radius
Theorem B.1.
Let $N_{a}(n)$ be the number of pulls for arm $a$ at round $n$ and $n=\sum_{a\in\mathcal{A}}N_{a}(n)$ is the total number of pulls. Let $\hat{\mu}_{a}(n)$ be the empirical mean of arm $a$ at round $n$ and $\hat{\sigma}_{a}^{2}(n)$ be the empirical variance of arm $a$ at round $n$. We define the empirical Bernstein Confidence Radius $\beta_{a}(n)$ as
$\beta_{a}(n)\;=\;\sqrt{\frac{2\,\hat{\sigma}_{a}^{2}\,\ln n}{\max(1,N_{a}(n))}}\;+\frac{3\,\ln n}{\max(1,N_{a}(n))}.$
Then we have the following bound holds with probability at least $1-\frac{6}{n}$:
$\left|\hat{\mu}_{a}-\mu_{a}\right|\leq\beta_{a}(n)$
Proof.
Under the setting of frame-query relevance setting, the reward $r_{t}$ and latent frame reward $y_{t}$ is naturally bounded in $[0,1]$. Therefore, according to Bernstein inequality, for any $\delta\in(0,1)$, we have
$\mathcal{P}\left[\mu_{a}\leq\hat{\mu}_{a}(n)+\sqrt{\frac{2\hat{\sigma}_{a}^{2}\ln\frac{3}{\delta}}{N_{a}(n)}}+\frac{3\ln\frac{3}{\delta}}{N_{a}(n)}\right]\geq 1-\delta.$
And symmetrically, we have
$\mathcal{P}\left[\mu_{a}\geq\hat{\mu}_{a}(n)-\sqrt{\frac{2\hat{\sigma}_{a}^{2}\ln\frac{3}{\delta}}{N_{a}(n)}}-\frac{3\ln\frac{3}{\delta}}{N_{a}(n)}\right]\geq 1-\delta.$
Therefore, we have
$\mathcal{P}\left[\left|\hat{\mu}_{a}-\mu_{a}\right|\leq\sqrt{\frac{2\hat{\sigma}_{a}^{2}\ln\frac{3}{\delta}}{N_{a}(n)}}+\frac{3\ln\frac{3}{\delta}}{N_{a}(n)}\right]\geq 1-2\delta.$
Choose $\delta=\frac{3}{n}$, then we have
$\left|\mu_{a}-\hat{\mu}_{a}(n)\right|\leq\sqrt{\frac{2\hat{\sigma}_{a}^{2}\ln\frac{3}{\delta}}{N_{a}(n)}}+\frac{3\ln\frac{3}{\delta}}{N_{a}(n)}.$
holds with probability at least $1-\frac{6}{n}$.
When $N_{a}(n)=0$, the statement is trivially true. Thus, we have the following bound holds with probability at least $1-\frac{6}{n}$:
$\left|\mu_{a}-\hat{\mu}_{a}(n)\right|\leq\beta_{a}(n).$
Appendix C Regret Bound
Arm-level Regret Bound
Theorem C.1.
Algorithm 2 returns the oracle top-$s$ set $S^{\star}$ with probability at least $1-\frac{6M}{n}$ when terminated.
Proof.
When Algorithm 2 terminates, the following condition holds:
$\max_{a\notin\hat{S}}\hat{\mu}_{n}(a)+\beta_{a}(n)\leq\min_{a\in\hat{S}}\hat{\mu}_{n}(a)-\beta_{a}(n).$
According to Theorem B.1, with probability at least $1-\frac{6}{n}$, we have $\left|\mu_{a}-\hat{\mu}_{a}(n)\right|\leq\beta_{a}(n)$ for all arms $a$. Therefore, for any $a\notin\hat{S}$,
$\mathcal{P}\left[a\in S^{\star}\right]\leq 1-\frac{6}{n}.$
Thus, the probability that there does not exist such an arm $a$ is at least $1-\frac{6(M-m)}{n}$, where $m$ is size of the $\hat{S}$ set. And this completes the proof. ∎
Frame-level Regret Bound
We define the frame-level regret as the difference between the optimal frame-level reward and the reward of the selected frames.
$r_{N}^{\mathrm{frame}}=\sum_{t\in{\mathbb{K}}^{\star}}y_{t}-\sum_{t\in\widehat{{\mathbb{K}}}_{n}}y_{t}.$
As long as we obtain the oracle top-$s$ set $S^{\star}$, the frame-level regret is also guaranteed to be small. As Frame-level sampling is actually finite so we can always find the top-$k$ frames with the highest rewards.
$\mathbb{E}r_{N}^{\mathrm{frame}}=\mathbb{E}\sum_{t\in{\mathbb{K}}^{\star}}y_{t}-\sum_{t\in\widehat{{\mathbb{K}}}_{n}}y_{t}\\ =\mathbb{E}\sum_{a\in S^{\star}}\sum_{t\in{\mathbb{K}}_{a}^{\star}}2\epsilon_{\psi}=0.$
For tighter bound, we leave this to future work.
Appendix D Limitations
In this work, we assume the frame-query relevance scores are i.i.d. and the temporal dependencies between frames are not considered. However, in practice, the frame-query relevance scores are dependent on the temporal dependencies between frames. As different parts may have strong correlations, this assumption may not hold. In this setting, we can use the Lipschitz/metric bandit problem (Kleinberg et al., 2008; Bubeck et al., 2011) or contextual bandit problem (Chu et al., 2011; Agarwal et al., 2014) to model the problem. We leave this as future work.
Appendix E The Use of Large Language Models (LLMs)
We used GPT-5 and Claude 4 solely for proofreading and light copy-editing (typos, grammar, and minor phrasing). All technical content, scientific claims, mathematical proofs, algorithms, experiment design and execution, dataset handling, figures, and evaluations were authored and verified by the human authors. LLMs were not used to generate ideas, code, data, results, or reviews; they did not contribute content at the level of a co-author. All suggested edits were manually inspected and accepted or rejected by the authors.